OpenStudy (anonymous):

I need some help please! Use the Binomial Theorem to expand the binomial. (3v + s)^5 a.) s5 – 5s4v + 10s3v2 – 10s2v3 + 5sv4 – v5 b.) s5 + 15s4v + 90s3v2 + 270s2v3 + 405sv4 + 243v5 c.) s5 + 45s4v + 270s3v2 + 810s2v3 + 1,215sv4 + 729v5 d.) s5 + 15s4 + 90s3 + 270s2 + 405s + 243

OpenStudy (anonymous):

there is not a minus sign in sight for forget the first answer.

OpenStudy (anonymous):

\((s+3v)^5\) first term is \(s^5\) second term is \(5\times s^4\times 3v\) or \(15s^4v\) the 5 because \(_5C_1=5\)

OpenStudy (anonymous):

that means C is out as well as A

OpenStudy (anonymous):

D is also out because it doesn't even have a \(v\) in it. so i guess by process of elimination it must be B

OpenStudy (anonymous):

Alright thanks :D

OpenStudy (anonymous):

i also cant find the fourth term of (d – 4b)^3 :/ think you can help?

OpenStudy (anonymous):

terms of \((a+b)^5\) look like \[a^5+5a^4b+10a^3b^2+10a^2b^3+5ab^4+b^5\] you can get the coefficients off of pascal's triangle

OpenStudy (anonymous):

heres the possible choice for it b3 –b3 64b3 –64b3

OpenStudy (anonymous):

there are only 4 terms in \((a-4b)^3\) the last term is \((-4b)^3=-64b^3\)

OpenStudy (anonymous):

alright, thanks a ton! :D

OpenStudy (anonymous):

yw